# From Understanding to Utilization: A Survey on Explainability for Large Language Models

### 23 Jan 2024 v1

@@ Abstract
This survey paper delves into the burgeoning field of explainability for Large Language Models (LLMs), a critical yet challenging aspect of natural language processing. With LLMs playing a pivotal role in various applications, their “black-box” nature raises concerns about transparency and ethical use. This paper emphasizes the necessity for enhanced explainability in LLMs, addressing both the general public’s trust and the technical community’s need for a deeper understanding of these models. We concentrate on pre-trained Transformer-based LLMs, such as LLaMA Touvron et al. (2023), which present unique interpretability challenges due to their scale and complexity. Our review categorizes existing explainability methods and discusses their application in improving model transparency and reliability. We also discuss representative evaluation methods, highlighting their strengths and limitations. The goal of this survey is to bridge the gap between theoretical understanding and practical application, offering insights for future research and development in the field of LLM explainability.
@@ 1 Introduction
In the rapidly evolving field of natural language processing, Large Language Models (LLMs) have emerged as a cornerstone, demonstrating remarkable proficiency across a spectrum of tasks. Despite their effectiveness, LLMs, often characterized as “black-box” systems, present a substantial challenge in terms of explainability and transparency. This opacity can lead to unintended consequences, such as generating harmful or misleading content Gehman et al. (2020), and the occurrence of model hallucinations Weidinger et al. (2021). These issues underscore the urgency for enhanced explainability, not just for understanding but for responsible and ethical application.
Explainability in LLMs serves two critical functions. For end users, it fosters trust by clarifying the model’s reasoning in a non-technical manner, enhancing understanding of their capabilities and potential flaws Zhao et al. (2023). For developers and researchers, it offers insights into unintended biases and areas for improvement, serving as a tool for refining model performance on downstream tasks Bastings et al. (2022); Meng et al. (2023a); Li et al. (2023b). However, the scale of LLMs poses unique challenges for explainability. Larger models with more parameters and extensive training data are harder to interpret. Traditional explanation methods like SHAP values Lundberg and Lee (2017) become less practical for these large-scale models Zhao et al. (2023). Moreover, a comprehensive understanding of LLM-specific phenomena, including in-context learning Halawi et al. (2023); Hendel et al. (2023); Todd et al. (2023); Wang et al. (2023), along with addressing issues like model hallucinations Ji et al. (2023); Chuang et al. (2023) and inherent biases dev (2023); An and Rudinger (2023); Schick et al. (2021), is vital for the ongoing refinement in model design.
In this literature review, we focus on explainability methods for pre-trained Transformer-based LLMs, often termed as base models. These models often scale up in training data and have billions of parameters, examples include GPT-2 Radford et al. (2019), GPT-J Chen et al. (2021), GPT-3 Brown et al. (2020), OPT Yordanov et al. (2022), and LLaMA family Touvron et al. (2023). In Section 2, we categorize and pose research questions based on our literature review. Based on this categorization, we review explainability methods in Section 3, followed by a discussion in Section 4 on how these insights are leveraged. We further discuss the evaluation methods and metrics in Section 5. Our goal is to synthesize and critically assess contemporary research, aiming to bridge the gap between understanding and practical application of insights derived from complex language models.
@@ 2 Overview
The field of large language models LLMs is rapidly advancing, making explainability not only a tool for understanding these complex systems but also essential for their improvement. This section categorizes current explainability approaches, highlights the challenges in ethical and controllable generation, and proposes research questions for future exploration.
@@ 3.1 Local Analysis
Local explanations in LLMs aim to elucidate how models generate specific predictions, such as sentiment classification or token predictions, for a given input. This section categorizes local explanation methods into two types: feature attribution analysis, and the analysis into individual Transformer Vaswani et al. (2017) components.
3.1.1 Feature Attribution Analysis
Feature attribution analysis quantifies the relevance of each input token to a model’s prediction. Given an input text $x$ with $n$ tokens $\{x_{1},x_{2},...,x_{n}\}$, a pre-trained language model $f$ outputs $f(x)$. Attribution methods assign a relevance score $R(x_{i})$ Modarressi et al. (2022); Ferrando et al. (2022); Modarressi et al. (2023) to each token $x_{i}$, reflecting its contribution to $f(x)$. This category includes perturbation-based, gradient-based, and vector-based methods.
@@ 3.1.1 Feature Attribution Analysis
One might consider gradient-based explanation methods as a natural approach for feature attribution. This type of method compute per-token importance scores Kindermans et al. (2016) using backward gradient vectors. Techniques like gradient $\times$ input Kindermans et al. (2017) and integrated gradients (IG) Sundararajan et al. (2017) accumulate the gradients obtained as the input is interpolated between a reference point and the actual input. Despite their widespread use, one main challenge of IG is the computational overhead to achieve high-quality integrals Sikdar et al. (2021); Enguehard (2023) Their attribution score has also shown to be unreliable in terms of faithfulness Ferrando et al. (2022) and their ability to elucidate the forward dynamics within hidden states remains constrained.
@@ 3.1.2 Dissecting Transformer Blocks
Tracking Transformer block’s component-by-component internal processing can provide rich information about their intermediate processing, given the stacked architecture of decoder-based language models Kobayashi et al. (2023). In a transformer inference pass, the input embeddings are transformed through a sequence of $L$ transformer layers, each composed of a multi-head self-attention (MHSA) sublayer followed by an MLP sublayer Vaswani et al. (2017). Formally, the representation $x_{i}^{l}$ of token $i$ at layer $l$ is obtained by:
@@ 3.2.2 Mechanistic Interpretability
Figure 3: The intensity of each grid cell represents the average causal indirect effect of a hidden state on expressing a factual association. Darker cells indicate stronger causal mediators. It was found that MLPs at the last subject token and attention modules at the last token play crucial roles. Meng et al. (2023a)
@@ 4.1 Model Editing
This type of knowledge editors include memory-based model and editors with additional parameters. Memory-based models store all edit examples explicitly in memory based on the explainability finding of key-value memories inside the FFN (Section 3.1.2). They can then employ a retriever to extract the most relevant edit facts for each new input, guiding the model to generate the edited fact. SERAC Mitchell et al. (2022), for instance, adopts a distinct counterfactual model while leaving the original model unchanged. Editors with additional parameters introduce extra trainable parameters within LLMs. These parameters are trained on a modified dataset while the original model parameters remain static. For example, T-Patcher Huang et al. (2023) integrates one neuron (patch) for one mistake in the last layer of the FFN of the model, which takes effect only when encountering its corresponding mistake.
@@
The locate-then-edit paradigm first identifies parameters corresponding to specific knowledge and then modifies them through direct updates to the target parameters. The Knowledge Neuron (KN) method Dai et al. (2022) introduces a knowledge attribution technique to pinpoint the “knowledge neuron” (a key-value pair in the FFN matrix) that embodies the knowledge and then updates these neurons. ROME Meng et al. (2023a) and MEMIT Meng et al. (2023b) apply causal tracing (Section 3.2.2) to locate the editing area. Instead of modifying the knowledge neurons in the FFN, ROME alters the entire matrix. Building upon these two methods, PMET Li et al. (2023c) involves the attention value to achieve better performance.
@@ 4.3.1 Reducing Hallucination
Hallucinations in LLMs refer to generated content not based on training data or facts, various factors like imperfect learning and decoding contribute to this Ji et al. (2023). To mitigate hallucinations, initial approaches used reinforcement learning from human feeback Ouyang et al. (2022) and distillation into smaller models like Alpaca Li et al. (2023d). Leveraging explainability provides a significantly less expensive way to reduce hallucination, enjoying the advantage of being adjustable and minimally invasive. For example, Li et al. (2023b) use as few as 40 samples to locate and find “truthful” heads and directions through a trained probe (Section 3.2.1). They propose inference-time intervention (ITI), a computationally inexpensive strategy to intervene on the attention head to shift the activations in the “truthful” direction, which achieves comparable or better performance towards instruction-finetuned model.
@@ 4.3.2 Ethical Alignment
As research on AI fairness gains increasing importance, there have been efforts to detect social bias Fleisig et al. (2023); An and Rudinger (2023) and suppress toxicity Gehman et al. (2020); Schick et al. (2021) in LMs. Many previous debiasing methods Qian et al. (2022) have focused on constructing anti-stereotypical datasets and then either retraining the LM from scratch or conducting fine-tuning. This line of debiasing approaches, although effective, comes with high costs for data construction and model retraining. Moreover, it faces the challenge of catastrophic forgetting if fine-tuning is performed Zhao et al. (2023). While few work has focused on the interpretability of the fairness research, dev (2023) explore interpreting and mitigating social biases in LLMs by introducing the concept of social bias neurons. Inspired by the gradient-based attribution method IG (Section 3.1.1), they introduce an interpretable technique, denoted as intergrated gap gradient (IG2), to pinpoint social bias neurons by back-propagating and integrating the gradients of the logits gap for a selected pair of demographics  [^3] Demographic include properties like gender, sexuality, occupation, etc. 9 common demographics are collected and pairs of demographics are selected to reveal the fairness gap dev (2023). Derived from this interpretation, they suppress the pinpointed neuron activations to mitigate the bias. Extensive experiments have verified the effectiveness of this method and have yielded the potential applicability of the explainability method for ethical alignment research in LLMs.
@@ 5 Evaluation
Recently, LLMs such as GPT-4 OpenAI (2023) have shown impressive abilities to generate natural language explanations for their predictions. However, it remains unclear whether these explanations actually assist humans in understanding the model’s reasoning Zhao et al. (2023). Specifically designed evaluation methods are needed to better assess the performance of explainability methods, such as attribution. Furthermore, calibrated datasets and metrics are required to evaluate the application of explainability to downstream tasks, such as truthfulness evaluation  [^4] Due to space limit, we only discusse the most commonly used evaluation approaches in explainability research.
@@ 5.1 Evaluating Explanation Plausibility
One common technique to evaluate the plausibility of attribution explanation is to remove Kof tokens with the highest or lowest estimated importance to observe its impact on the model’s output Chen et al. (2020); Modarressi et al. (2023). Another approach to assessing explanation plausibility involves indirect methods, such as measuring the performance of model editing, particularly for “locate-then-edit” editing methods, which heavily rely on interpretation accuracy. Recent research Yao et al. (2023); Zhao et al. (2023) suggests that having evaluation datasets is crucial for evaluating factual editing in LLMs. Two commonly used datasets for this purpose are ZsRE Levy et al. (2017), a Question Answering (QA) dataset that employs question rephrasings generated through back-translation, and CounterFact Meng et al. (2023a), a more challenging dataset that includes counterfacts starting with low scores in comparison to correct facts.
@@ 6 Conclusion
In this paper, we have presented a comprehensive overview on explainability for LLMs and their applications. We have summarized methods for local and global analysis based on the objectives of explanations. Additionally, we have discussed the use of explanations to enhance models and the evaluation of these methods. Major future research directions in understanding LLMs include developing explanation methods tailored to different language models and making LLMs more trustworthy and aligned with human values by leveraging explainability knowledge. As LLMs continue to advance, explainability will become incredibly vital to ensure that these models are transparent, fair, and beneficial. We hope that this literature review provides a useful overview of this emerging research area and highlights open problems and directions for future research.

### 22 Feb 2024 v2

@@ Abstract
Explainability for Large Language Models (LLMs) is a critical yet challenging aspect of natural language processing. As LLMs are increasingly integral to diverse applications, their “black-box” nature sparks significant concerns regarding transparency and ethical use. This survey underscores the imperative for increased explainability in LLMs, delving into both the research on explainability and the various methodologies and tasks that utilize an understanding of these models. Our focus is primarily on pre-trained Transformer-based LLMs, such as LLaMA Touvron et al. (2023), which pose distinctive interpretability challenges due to their scale and complexity. In terms of existing methods, we classify them into local and global analyses, based on their explanatory objectives. When considering the utilization of explainability, we explore several compelling methods that concentrate on model editing, control generation, and model enhancement. Additionally, we examine representative evaluation metrics and datasets, elucidating their advantages and limitations. Our goal is to reconcile theoretical and empirical understanding with practical implementation, proposing exciting avenues for explanatory techniques and their applications in the LLMs era.
@@ 1 Introduction
In the rapidly evolving field of natural language processing, Large Language Models (LLMs) have emerged as a cornerstone, demonstrating remarkable proficiency across a spectrum of tasks. Despite their effectiveness, LLMs, often characterized as “black-box” systems, present a substantial challenge in terms of explainability and transparency. This opacity can lead to unintended consequences, such as the generation of harmful or misleading content Gehman et al. (2020), and the occurrence of model hallucinations Weidinger et al. (2021). These issues underscore the urgency for improved explainability, not just for understanding, but for responsible and ethical application.
Explainability in LLMs serves two critical functions. For end users, it fosters trust by clarifying the model’s reasoning in a nontechnical manner, enhancing understanding of their capabilities and potential flaws Zhao et al. (2023). For developers and researchers, it offers insights into unintended biases and areas of improvement, serving as a tool for improving the performance of the model in downstream tasks Bastings et al. (2022); Meng et al. (2023a); Li et al. (2023b). However, the scale of LLMs poses unique challenges to explainability. Larger models with more parameters and extensive training data are harder to interpret. Traditional explanation methods such as SHAP values Lundberg and Lee (2017) become less practical for these large-scale models Zhao et al. (2023). Moreover, a comprehensive understanding of LLM-specific phenomena, including in-context learning Halawi et al. (2023); Hendel et al. (2023); Todd et al. (2023); Wang et al. (2023), along with addressing issues such as model hallucinations Ji et al. (2023); Chuang et al. (2023) and inherent biases dev (2023); An and Rudinger (2023); Schick et al. (2021), is vital for ongoing refinement in model design.
In this survey, we focus on explainability methods for pre-trained Transformer-based LLMs, often termed as base models. These models often scale up in training data and have billions of parameters; examples include GPT-2 Radford et al. (2019), GPT-J Chen et al. (2021), GPT-3 Brown et al. (2020), OPT Yordanov et al. (2022), and LLaMA family Touvron et al. (2023). In Section 2, we categorize and pose research questions based on our survey. Based on this categorization, we review explainability methods in Section 3, followed by a discussion in Section 4 on how these insights are leveraged. We further discuss the evaluation methods and metrics in Section 5. Our goal is to synthesize and critically assess contemporary research, aiming to bridge the gap between understanding and practical application of insights derived from complex language models.
@@ 2 Overview
The field of LLMs is rapidly advancing, making explainability not only a tool for understanding these complex systems but also essential for their improvement. This section categorizes current explainability approaches, highlights the challenges in ethical and controllable generation, and proposes research questions for future exploration.
@@ 3.1 Local Analysis
Local explanations in LLMs aim to elucidate how models generate specific predictions, such as sentiment classification or token predictions, for a given input. This section categorizes local explanation methods into two types: feature attribution analysis and analysis into individual Transformer Vaswani et al. (2017) components.
3.1.1 Feature Attribution Explanation
Feature attribution, a local method for explaining a prediction, analysis quantifies the relevance of each input token to a model’s prediction. Given an input text $x$ with $n$ tokens $\{x_{1},x_{2},...,x_{n}\}$, a pre-trained language model $f$ outputs $f(x)$. Attribution methods assign a relevance score $R(x_{i})$ Modarressi et al. (2022); Ferrando et al. (2022); Modarressi et al. (2023) to each token $x_{i}$, reflecting its contribution to $f(x)$. This category includes perturbation-based, gradient-based, and vector-based methods.
@@ 3.1.1 Feature Attribution Explanation
One might consider gradient-based explanation methods as a natural approach for feature attribution. This type of method computes per-token importance scores Kindermans et al. (2016) using backward gradient vectors. Techniques such as gradient $\times$ input Kindermans et al. (2017) and integrated gradients (IG) Sundararajan et al. (2017) accumulate the gradients obtained as the input is interpolated between a reference point and the actual input. Despite their widespread use, one main challenge of IG is the computational overhead to achieve high-quality integrals Sikdar et al. (2021); Enguehard (2023) Their attribution score has also shown to be unreliable in terms of faithfulness Ferrando et al. (2022) and their ability to elucidate the forward dynamics within hidden states remains constrained.
@@ 3.1.2 Dissecting Transformer Blocks
Tracking Transformer block’s component-by-component internal processing can provide rich information on its intermediate processing, given the stacked architecture of decoder-based language models Kobayashi et al. (2023). In a transformer inference pass, the input embeddings are transformed through a sequence of $L$ transformer layers, each composed of a multi-head self-attention (MHSA) sublayer followed by an MLP sublayer Vaswani et al. (2017). Formally, the representation $x_{i}^{l}$ of token $i$ at layer $l$ is obtained by:
@@ 3.2.2 Mechanistic Interpretability
Figure 3: The intensity of each grid cell represents the average causal indirect effect of a hidden state on expressing a factual association. Darker cells indicate stronger causal mediators. It was found that the MLPs at the last subject token and the attention modules at the last token play crucial roles. Meng et al. (2023a)
@@ 4.1 Model Editing
This type of knowledge editors includes memory-based model and editors with additional parameters. Memory-based models store all edit examples explicitly in memory based on the explainability finding of key-value memories inside the FFN (Section 3.1.2). They can then employ a retriever to extract the most relevant edit facts for each new input, guiding the model to generate the edited fact. SERAC Mitchell et al. (2022), for instance, adopts a distinct counterfactual model while leaving the original model unchanged. Editors with additional parameters introduce extra trainable parameters within LLMs. These parameters are trained on a modified dataset while the original model parameters remain static. For example, T-Patcher Huang et al. (2023) integrates one neuron (patch) for one mistake in the last layer of the FFN of the model, which takes effect only when encountering its corresponding mistake.
@@
The locate-then-edit paradigm first identifies the parameters corresponding to the specific knowledge and then modifies them by directly updating the target parameters. The Knowledge Neuron (KN) method Dai et al. (2022) introduces a knowledge attribution technique to pinpoint the “knowledge neuron” (a key-value pair in the FFN matrix) that embodies the knowledge and then updates these neurons. ROME Meng et al. (2023a) and MEMIT Meng et al. (2023b) apply causal tracing (Section 3.2.2) to locate the editing area. Instead of modifying the knowledge neurons in the FFN, ROME alters the entire matrix. Based on these two methods, PMET Li et al. (2023c) involves the attention value to achieve better performance.
@@ 4.3.1 Reducing Hallucination
Hallucinations in LLMs refer to generated content not based on training data or facts, various factors such as imperfect learning and decoding contribute to this Ji et al. (2023). To mitigate hallucinations, initial approaches used reinforcement learning from human feeback Ouyang et al. (2022) and distillation into smaller models such as Alpaca Li et al. (2023d). Leveraging explainability provides a significantly less expensive way to reduce hallucination, enjoying the advantage of being adjustable and minimally invasive. For example, Li et al. (2023b) use as few as 40 samples to locate and find “truthful” heads and directions through a trained probe (Section 3.2.1). They propose inference-time intervention (ITI), a computationally inexpensive strategy to intervene on the attention head to shift the activations in the “truthful” direction, which achieves comparable or better performance toward the instruction-finetuned model.
@@ 4.3.2 Ethical Alignment
As research on AI fairness gains increasing importance, there have been efforts to detect social bias Fleisig et al. (2023); An and Rudinger (2023) and suppress toxicity Gehman et al. (2020); Schick et al. (2021) in LMs. Many previous debiasing methods Qian et al. (2022) have focused on constructing anti-stereotypical datasets and then either retraining the LM from scratch or conducting fine-tuning. This line of debiasing approaches, although effective, comes with high costs for data construction and model retraining. Moreover, it faces the challenge of catastrophic forgetting if fine-tuning is performed Zhao et al. (2023). While few work has focused on the interpretability of the fairness research, dev (2023) explore interpreting and mitigating social biases in LLMs by introducing the concept of social bias neurons. Inspired by the gradient-based attribution method IG (Section 3.1.1), they introduce an interpretable technique, denoted as intergrated gap gradient (IG2), to pinpoint social bias neurons by back-propagating and integrating the gradients of the logits gap for a selected pair of demographics  [^3] Demographic include properties like gender, sexuality, occupation, etc. 9 common demographics are collected and pairs of demographics are selected to reveal the fairness gap dev (2023). Taking this interpretation, they suppress the activations of the pinpointed neurons to mitigate bias. Extensive experiments have verified the effectiveness of this method and have yielded the potential applicability of the explainability method for ethical alignment research in LLMs.
@@ 5 Evaluation
Recently, LLMs such as GPT-4 OpenAI (2023) have shown impressive abilities to generate natural language explanations for their predictions. However, it remains unclear whether these explanations actually help humans understand the reasoning of the model Zhao et al. (2023). Specifically designed evaluation methods are needed to better assess the performance of explainability methods, such as attribution. Furthermore, calibrated datasets and metrics are required to evaluate the application of explainability to downstream tasks, such as truthfulness evaluation  [^4] Due to space limit, we only discusse the most commonly used evaluation approaches in explainability research.
@@ 5.1 Evaluating Explanation Plausibility
One common technique to evaluate the plausibility of attribution analysis is to remove Kof tokens with the highest or lowest estimated importance to observe its impact on the model output Chen et al. (2020); Modarressi et al. (2023). Another approach to assessing explanation plausibility involves indirect methods, such as measuring the performance of model editing, particularly for “locate-then-edit” editing methods, which heavily rely on interpretation accuracy. Recent research Yao et al. (2023); Zhao et al. (2023) suggests that having evaluation datasets is crucial for evaluating factual editing in LLMs. Two commonly used datasets for this purpose are ZsRE Levy et al. (2017), a Question Answering (QA) dataset that employs question rephrasings generated through back-translation, and CounterFact Meng et al. (2023a), a more challenging dataset that includes counterfacts starting with low scores compared to correct facts.
@@ 6 Conclusion
In this survey, we have presented a comprehensive overview of explainability for LLMs and their applications. We have summarized methods for local and global analysis based on the objectives of explanations. In addition, we have discussed the use of explanations to enhance models and the evaluation of these methods. Major future research directions to understanding LLM include developing explanation methods tailored to different language models and making LLMs more trustworthy and aligned with human values by using explainability knowledge. As LLMs continue to advance, explainability will become incredibly vital to ensure that these models are transparent, fair, and beneficial. We hope that this review of the literature provides a useful overview of this emerging research area and highlights open problems and directions for future research.